OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x^2 − 16x + ____

OpenStudy (anonymous):

what is half of 16?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

what is 8 squared?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

you win

OpenStudy (anonymous):

what? is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok thank you :)

OpenStudy (anonymous):

no, i knew it already

OpenStudy (anonymous):

\[x^2-16x+64=(x-8)(x-8)=(x-8)^2\]

OpenStudy (skullpatrol):

@babyblue123 Did you learn anything?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

What?

OpenStudy (anonymous):

oh ok thank you :)

OpenStudy (anonymous):

what? is that the answer?

OpenStudy (anonymous):

64

OpenStudy (skullpatrol):

What did you learn?